ESCAPING THE FLOOD OF TIME: NOAH’S ARK IN W. G. SEBALD’S AUSTERLITZ

CARL S. MCTAGUE

On 14 December 2001, German expatriate writer W. G. Sebald died in a car accident in East Anglia. He had been at the hight of his literary career. The English translations of his novels The Emigrants (1997) and The Rings of Saturn (1999) had only recently established his position among his generation’s best writers—both German and English. Only a few months before his untimely death, a translation of what would be his final novel, Austerlitz, was published. Through its unnamed narrator, the reader follows Jacques Austerlitz’s personal struggle to uncover his deeply repressed identity as a Czech refugee who as a young child fled Nazi territory on the Kindertransport.

In this essay I will examine what might at first appear to be a very narrow topic within Austerlitz: two appearances of Noah’s Ark. The first is a golden picture of the Ark (reproduced on page forty-three of the text) in the Freemasons’ temple of the Great Eastern Hotel in London. It is while gazing at this “ornamental gold-painted picture of a three-story ark floating beneath a rainbow, and the dove just returning to it carrying the olive branch in her beak” (43) that Austerlitz decides that “he must find someone to whom he could relate his own story, a story which he had learned only in the last few years and for which he needed the kind of listener I [the narrator] had once been in Antwerp, Liège, and Zeebrugge” (43-44).

The second Ark appears much later, when Austerlitz describes visiting the hermetically sealed billiards room of Iver Grove, an abandoned estate on the outskirts of postwar Oxford. Therein, “sealed away so long from the flow of the hours and days and the succession of the generations” (108), was a toy “Noah’s Ark with the pairs of well-behaved animals saved from the Flood looking out of it” (108).

I will interpret these appearances by connecting the Ark, and the devastating flood from which it enables Noah to escape, to a broad and pervasive theme of the novel: the metaphor of water as time. Austerlitz himself makes this metaphor explicit in a conversation with the narrator at the Royal Observatory just after photographing a pocket watch:

If Newton thought, said Austerlitz, pointing through the window and down to the curve of the water around the Isle of Dogs glistening in the last of the daylight, if Newton really thought that time was a river like the Thames, then where is its source and into what sea does it finally flow? Every river, as we know, must have banks on both sides, so where, seen in those terms, where are the banks of time? What would be this river’s qualities, qualities perhaps corresponding to those of water, which is fluid, rather heavy, and translucent? In what way do objects immersed in time differ from those left untouched by it? Why do we show the hours of light and darkness in the same circle? Why does time stand eternally still and motionless in one place, and rush headlong by in another? Could we not claim, said Austerlitz, that time itself has been nonconcurrent over the centuries and the millennia? It is not so long ago, after all, that it began spreading out over everything. And is not human life in many parts of the earth governed to this day less by time than by the weather, and thus by an unquantifiable dimension which disregards linear regularity, does not progress constantly forward but moves in eddies, is marked by episodes of congestion and irruption, recurs in ever-changing form, and evolves in no one knows what direction? (100-101)

Answers to Austerlitz’s questions about time, especially those concerning the global geometry of time’s flow and time’s effect on submerged objects, are central to a comprehensive interpretation of Sebald’s water-time metaphor, and hence also to that of the Ark references. Unfortunately, precise answers to these questions still elude my grasp. I will therefore discuss the ambiguous and conflicting clues which currently obstruct my interpretation. From these clues, the coarse strokes of an incomplete solution will emerge, a solution which fortunately nevertheless suffices to interpret the narrower question of the Ark’s function.

Perhaps the novel’s earliest and most concrete image of submersion is Austerlitz’s childhood fascination with his foster father Elias’s home village, Llanwddyn, which had been abandoned and submerged when the Vyrnwy dam was built in 1888. Inspired by Elias’s album of family photographs, young Austerlitz vividly pictured the inhabitants of Llanwddyn still alive in the reservoir’s depths, and even imagined himself among these silent, ghostly figures:

I often felt as if I too had been submerged in that dark water, and like the poor souls of Vyrnwy must keep my eyes wide open to catch a faint glimmer of light far above me, and see the reflection, broken by ripples, of the stone tower standing in such fearsome isolation on the wooded bank. (53)^{1}

Because the inhabitants of Llanwddyn had not perished in the flood, and were in fact almost certainly, like Elias himself, still alive at the time of Austerlitz’s visions, the Llanwddyn passage strongly suggests that submersion to Sebald is connected not directly with death, but instead with the loss of memory: people and objects submerged in time are forgotten. Young Austerlitz’s imaginary act of diving into the depths of the Vyrnwy reservoir signifies his attempt to remember the forgotten village.

1Note that this reference to the stone tower seems to suggest a link with Austerlitz’s recollections of the so-called Mäuseturm along the part of the Rhine known as Bringer Loch (225), a passage which seems, at least cryptically, to address the preceding water-time questions, including, “where are the banks of time?” (101).

This interpretation of submersion is supported by several other passages. For instance, Austerlitz tells the narrator:

In my photographic work I was always especially entranced, said Austerlitz, by the moment when the shadows of reality, so to speak, emerge out of nothing on the exposed paper, as memories do in the middle of the night, darkening again if you try to cling to them, just like a photographic print left in the developing bath too long. (77)

Closely linked to submersion is the recurring image of the snow-covered ground. The first is the narrator’s “childhood wish for everything to be snowed over, my [the narrator’s] whole village and the valley all the way to the mountain peaks, and how I used to imagine what it would be like when we thawed out again and emerged from the ice in spring” (37). Even less ambiguous is Austerlitz’s childhood fascination with “winter pictures […] showing hares, deer, and partridges transfixed with astonishment as they stared at the ground covered with newly fallen snow” (204). These images are strongly connected to memory by the question which “constantly troubled” (204) young Austerlitz: “if it’s all white, how do the squirrels know where they’ve buried their hoard?” (204). Austerlitz’s nanny Vera emphasizes the broader significance of his childhood concern when she asks the adult Austerlitz: “how indeed do the squirrels know, what do we know ourselves, how do we remember, and what is it we find in the end?” (204).

These passages draw our attention to Austerlitz’s fascination with memory, and his devotion to its preservation. In fact, the book could not altogether inaccurately be summarized as Austerlitz’s personal struggle to remember the past. This viewpoint leads to an interesting interpretation of the following cryptic passage, taken from Austerlitz’s visit to Theresienstadt, the concentration camp where, as he eventually learns, his mother had once been detained. Peering at objects in the window display of the town’s Antikos Bazar, Austerlitz ponders:

What secret lay behind […] the endless landscape painted round a lampshade in fine brushstrokes, showing a river running quietly through perhaps Bohemia or perhaps Brazil? And then there was the stuffed squirrel, already moth-eaten here and there, perched on the stump of a branch in a showcase the size of a shoebox, which had its beady button eye implacably fixed on me, and whose Czech name—veverka—I now recalled like the name of a longlost friend. What, I asked myself, said Austerlitz, might be the significance of the river never rising from any source, never flowing out into any sea but always back into itself, what was the meaning of veverka, the squirrel forever perched in the same position? (195-197)

Could it be that this squirrel represents Austerlitz himself, and that these cryptic comments are an abstract means of self reflection? Is Austerlitz here asking himself the purpose of his endless journey into the past?

The river flowing in circles “back into itself” (196) on the lampshade’s cylindrical surface clearly refers to Austerlitz’s question “where is [time’s] source and into what sea does it finally flow” (100), and its placement in the window display strongly suggests that the Antikos Bazar is intended to contribute to the larger water-time metaphor. In fact, seen as a vessel of objects from the past, the Antikos Bazar is not altogether unlike the Ark. But it is not an uplifting image. Its artifacts, although physically preserved, are lifeless, detached from their original associations, and superposed in an almost senseless amalgam. Because their meanings have been forgotten, they are, to Austerlitz, like the survivors of Llanwddyn, as good as lost.

One of the most powerful and abstract passages contributing to the watertime metaphor is Austerlitz’s cold-war visit to Marienbad, the well-known spa resort which Austerlitz’s family had visited just before World War II. He accompanies Marie, the woman he will later regret having lost through his own strange inability to live in the present. Of the reception clerk at the hotel Austerlitz recalls “that although he could not have been much over forty his forehead was wrinkled in fan-like folds above the root of his nose” (208) and that he “went through the necessary formalities without another word, very slowly, almost as if he were moving in a denser atmosphere than ours” (208). These are the first of several details which portray Marienbad as a resort submerged, for a man submerged in water would have wrinkled skin, be unable to speak, and move slowly. Moreover, there is a layer of dust on the desk of their hotel room, a detail which perhaps refers to the billiard table without dust in the sealed room which contained the toy Ark. Austerlitz initially felt deeply content lying next to Marie, but as he fell asleep, his “mind became gradually submerged” (211), and in the morning he “sat up and, like a man seasick, had to perch on the edge of the bed” (211). Looking out the window he sees “the grand hotels ranged in a semicircle rising to the heights, the Pacifik, the Atlantic, the Metropole, the Polonia and Bohemia with their rows of balconies, their corner turrets and roof ridges emerging from the morning mist like oceangoing steamers from a dark sea” (212). Walking through the deserted town, he “kept feeling as if someone else were walking beside [him]” (212).

Sebald thus intimately links this passage to Austerlitz’s childhood vision of the submerged village of Llanwddyn. Compare, for instance, Marie’s plea to Austerlitz: “why do I see your lips opening as if you were about to say something, maybe even cry out loud, and then I hear not the slightest sound?” (215-216) to Austerlitz’s childhood vision of the former inhabitants of Llanwddyn “still down in the depths, sitting in their houses and walking along the road, but unable to speak and with their eyes opened far too wide” (51-52). Austerlitz is thus metaphorically submerged in Marienbad’s past, in particular, the Marienbad he once visited with Vera and his parents.

When Marie confronts Austerlitz about remaining “unapproachable” (215) and “like a pool of frozen water” (215) (this after warning him about Schumann’s descent into insanity, leading to his attempted suicide—by leaping into the icy waters of the Rhine), the two are standing in a spa’s pump room. “It isn’t true that we need absence and loneliness” (216) she tells him; she is here trying to rescue him from his submersion in the past, and this is why Sebald sets the scene in the pump room.

With the meaning of submersion better understood, the functions of the two Ark references with which I initiated this essay become more comprehensible. The Ark in the Freemasons’ temple reminds Austerlitz that his unique and painful story will be forgotten unless he shares it with an attentive listener; to preserve his memories, he must place them in a vessel—the narrator—so that they are not consumed by the deluge of time. The toy Ark in the sealed billiard room of Iver Grove acts as a symbol for the room in which it is found, for although the flow of time has reduced Iver Grove to a storehouse for sacks of potatoes and grain, the hermetically sealed billiard room and its contents have been perfectly preserved for more than 150 years.

Finally, I examine what is perhaps the novel’s most subtle, and powerful, reference to the Ark. Austerlitz’s investigation into his father’s disappearance from Paris ends in frustration at the new Bibliothèque Nationale, the state-of-the-art library whose notoriously elaborate but unreliable automatic retrieval system has been widely reported. Its architecture (depicted in a photograph on page 279 of the text) is somewhat unusual; it consists of four twenty-two story towers set at the corners of a rectangular platform which encloses an immense tree-filled atrium. That we should consider this unusual structure as a ship is clear from Austerlitz’s description of its “deck”:

You might think, especially on days when the wind drives rain over this totally exposed platform, as it quite often does, said Austerlitz, that by some mistake you had found your way to the deck of the Berengaria or one of the other oceangoing giants, and you would be not in the least surprised if, to the sound of a wailing foghorn, the horizon of the city of Paris suddenly began rising and falling against the gauge of the towers as the great steamer pounded onwards through mountainous waves, or if one of the tiny figures, having unwisely ventured on deck, were swept over the rail by a gust of wind and carried far out into the wastes of the Atlantic waters. (277-278)

But Sebald links the Bibliothèque to the biblical Ark in particular in two subtle ways. First, there is the “inner courtyard and the curious nature reserve cut, so to speak, from the surface of the promenade deck and sunk two or three stories deep, which has been planted with about a hundred full-grown stone pines from the Forêt de Bord transported, how I do not know, to this place of banishment” (280) and the “two mythical squirrels said to have been brought to the library in the hope that they will increase and multiply, founding a large colony of their species in this artificial pine grove” (281). These squirrels refer not only to the pairs of animals Noah stowed in his Ark, but to Austerlitz’s childhood fascination with veverka, and the stuffed squirrel in the display window of Theresienstadt’s Antikos Bazar. Second, there are Austerlitz’s many comparisons of the library’s architecture to that of a Mesopotamian ziggurat; such comparisons help link the library to the Ark’s historical period.

But the library is clearly dysfunctional as an Ark, as Sebald establishes in several ways. For instance, instead of carrying olive branches back to the Ark in their beaks, “birds which had lost their way in the library forest flew into the mirror images of the trees in the reading room windows, struck the glass with a dull thud, and fell lifeless to the ground” (281).

More directly, although the library “according to one of the loathsome phrases now current is supposed to serve as the treasure house of our entire literary heritage” (282), it “proved useless in my [Austerlitz’s] search for any traces of my father who had disappeared from Paris more than fifty years ago” (282). Austerlitz moreover criticizes the library as “unwelcoming if not inimical to human beings” (276) and as “an utter absurdity, something that must have been devised […] to instill a sense of insecurity and humiliation in the poor readers” (278). Lemoine, a librarian who recognizes Austerlitz from the old library, believes that the new building “in both its entire layout and its near-ludicrous internal regulation seeks to exclude the reader as a potential enemy” (286), and “might be described […] as the official manifestation of the increasingly importunate urge to break with everything which still has some living connection to the past” (286). He moreover links the new building with “the dissolution, in line with the inexorable spread of processed data, of our capacity to remember” (286). In this way, the library might in fact be considered an ‘anti-Ark’.

This view of the library as an anti-Ark is reinforced by a revelation about the ground on which it is built, for Austerlitz recalled that when Lemoin took him to the eighteenth floor of one of the library’s four towers:

As soon as you looked down at the light-colored promenade deck and the darker crowns of the trees emerging from it, the pull exerted by the abyss below took hold of you, forcing you to step back. Sometimes, so Lemoine told me, said Austerlitz, he felt the current of time streaming round his temples and brow when he was up here, but perhaps, he added, that is only a reflex of the awareness formed in my mind over the years of the various layers which have been superimposed on each other to form the carapace of the city. Thus, on the waste land between the marshaling yard of the gare d’Austerlitz and the pont de Tolbiac where this Babylonian library now rises, there stood until the end of the war an extensive warehousing complex to which the Germans brought all the loot they had taken from the homes of the Jews of Paris. (287-288)

So massive was this highly organized operation of expropriation that it required the “entire pantechnicon fleet of the Paris Union of Furniture Removers and an army of no fewer than fifteen hundred removal men” (288) months to ruthlessly clear some forty-thousand apartments. The warehouse’s holdings were so vast that over “five hundred art historians, antique dealers, restorers, joiners, clockmakers, furriers, and couturiers” (289) were required to “put the goods coming into the depot in proper order and sort them by value and kind” (289). In fact, “there were even special cardboard cartons set aside to hold the rosin removed, for the sake of greater cleanliness, from confiscated violin cases” (289).

This entire warehouse complex, many of whose holdings are still unaccounted for, “is burried in the most literal sense beneath the foundations of our pharaonic President’s Grande Bilbiothèque” (289). Thus, the flood of time has already eroded the warehouse, and although the library’s exalted pretense is to preserve its people’s literary heritage, it has in fact physically supplanted an object which Austerlitz believes a true Ark should have helped preserve.

Near the beginning of the novel, Austerlitz states to the narrator what I view as the novel’s central theme:

The darkness does not lift but becomes yet heavier as I think how little we can hold in mind, how everything is constantly lapsing into oblivion with every extinguished life, how the world is, as it were, draining itself, in that the history of countless places and objects which themselves have no power of memory is never heard, never described or passed on. (24)

It is perhaps not altogether unreasonable to view Sebald’s novel, and the other stunning works of his opus, as his intended contribution to an Ark, his attempt to preserve not merely abstracted artifacts of the past, but the memories of people and places whose tragedies are in our own age lapsing into oblivion.

Works Cited

Sebald, W. G. Austerlitz. Trans. Anthea Bell. New York: Random House, 2001.

Acknowledgements

This essay would not exist were it not for the insightful suggestions, and encouragements, of UC Professor and literata Maria Romagnoli Brackett.